Consider a relation R=\ (1,1),(1, ), (0,2), (2,3) (3,1)) on the set A=\ 1,2,3\ Find transitive closure of the relation R using algorithm Warshall's

1
Expert's answer
2021-12-17T07:00:54-0500

Consider a relation "R=\{ (1,1),(1, 0), (0,2), (2,3), (3,1)\}" on the set "A=\{0, 1,2,3\}".


Let us find transitive closure of the relation "R" using Warshall's algorithm:


"W^{(0)}=M_R
=\begin{pmatrix}
0 & 0 & 1 & 0\\
1 & 1 & 0 & 0\\
0 & 0 & 0 & 1\\
0 & 1 & 0 & 0
\end{pmatrix}"



"W^{(1)}
=\begin{pmatrix}
0 & 0 & 1 & 0\\
1 & 1 & 1 & 0\\
0 & 0 & 0 & 1\\
0 & 1 & 0 & 0
\end{pmatrix}"


"W^{(2)}
=\begin{pmatrix}
0 & 0 & 1 & 0\\
1 & 1 & 1 & 0\\
0 & 0 & 0 & 1\\
1 & 1 & 1 & 0
\end{pmatrix}"


"W^{(3)}
=\begin{pmatrix}
0 & 0 & 1 & 1\\
1 & 1 & 1 & 1\\
0 & 0 & 0 & 1\\
1 & 1 & 1 & 1
\end{pmatrix}"


"M_{R^*}=W^{(4)}
=\begin{pmatrix}
1 & 1 & 1 & 1\\
1 & 1 & 1 & 1\\
1 & 1 & 1 & 1\\
1 & 1 & 1 & 1
\end{pmatrix}"


It follows that "R^*=A\times A" is a universal relation on the set "A."